# Guided Query Refinement: Multimodal Hybrid Retrieval with Test-Time Optimization

### v1

@@ 3.2 Results - Performance
Table 2: Percentage gain, in $\mathrm{NDCG}@5$, of hybrid retrieval over the primary retriever for ViDoRe 2. Each cell depicts average gain over 99 retriever pairs (33 multimodal base retrievers $\times$ 33 text retrievers).
Method Avg Biomed Lectures Economics ESG Human ESG Full
@@
Table 2 depicts the aggregated performance of the different hybrid retrieval methods on ViDoRe 2, presented as the average percentage gain relative to the base retriever over all pairs (Table 4 in the appendix lists the average absolute values). The ranking aggregation methods (RRF, Average Ranking) generally lead to a deterioration in performance. All aggregation methods benefit from parameter tuning, with tuned score aggregation methods achieving consistent gains over the base retriever. GQR outperforms all other hybrid retrieval variants, with an average gain of $3.9\%$ over the base multimodal retriever.
3.3 Results - Time and Memory
@@
To characterize the test-time costs of GQR and ColPali-based retrievers, we ran latency measurements on a single NVIDIA A100 GPU, measuring document retrieval for randomly sampled queries from each ViDoRe 2 subset (averaging across $100$ runs). Figure 2 depicts the quality-latency trade-offs of the different retrievers, where each point on the graph represents either a base retriever or a retriever with GQR (Appendix E shows full values). The plot illustrates the substantial costs of the strongest base model – Llama-Nemo (orange square, right), which attains $\mathrm{NDCG}@5=62.9$ at a cost of $2{,}591\,\text{ms}$ per query. Notably, our Colnomic GQR hybrid, with Linq as the refinement model (green diamond, left), reaches $\mathrm{NDCG}@5=62.7$ at $181\,\text{ms}$ ($\approx\!14\times$ faster), and the Colnomic hybrid with Jina (green circle) attains $\mathrm{NDCG}@5=63.0$ at $350\,\text{ms}$ ($\approx\!7\times$ faster), surpassing Nemo. Across base models, applying GQR increases latency by small relative measures ($\approx\!60-80$ ms with refinement from dense retrievers and $\approx\!200-350$ with refinement from a multi-vector retriever) with large gains in performance, shifting the Pareto frontier left and upwards.
Storage.
Figure 2: Latency–quality tradeoff in online querying. The $x$ axis is runtime in milliseconds for a single query, on a log scale, and the $y$ axis is the average evaluation score (NDCG@5). Marker color encodes the primary retriever; marker shape encodes the GQR complementary retriever, with squares indicating the primary retriever alone (without applying GQR).
@@ 4.1 Reranker Comparison
Cross-encoder rerankers are widely used in retrieval pipelines, applying full query to document interaction via self-attention to the top-K documents. Similarly to GQR, rerankers operate at test time on a candidate pool returned by a bi-encoder, and thus provide a natural point of comparison. We evaluate GQR against lightonai/MonoQwen2-VL-v0.1, an open-weights multimodal reranker (Chaffin & Lac, 2024). Figure 3 illustrates the latency-performance characteristics of GQR against reranking the top-$5$ (on the left) and the top-$10$ (on the right) candidates returned by each retriever. GQR is run with a default $K=10$ configuration and with Linq-Embed as the refinement model. The left plot illustrates that GQR outperforms a top-$5$ reranking pipeline on both the latency and performance axes. Against the top-$10$ reranking pipelines (right), GQR achieves close performance while being $21\times$ faster for Colnomic, $16\times$ faster for Jina (vision), and $2\times$ faster for Nemo. Across both comparisons, GQR remains on the Pareto front and indicates attractive latency performance trade-offs. We report the full results in Table 9.
@@ 4.2 GQR Design Choices
Prior works on test-time query optimization often run a second index search with the optimized query $z^{(T)}$, aiming to retrieve documents beyond the initial top-$K$ and increase recall (Sung et al., 2023; Gangi Reddy et al., 2025). This extra pass adds latency as it repeats index traversal and candidate generation. We test this modification to GQR where we perform an additional search with the optimized query over the full index, and observe no improvement in performance (Table 14). This suggests that the effects of GQR in our setting are largely confined to the original pool $\mathcal{C}(q)$ of candidate documents.
Candidate pool policy.
@@
Primary and complementary roles
@@ 5 Related Works
Recent advances in visual document retrieval have been dominated by late-interaction, multi-vector architectures. This paradigm, first introduced to the multimodal domain by ColPali (Faysse et al., 2025), adapts the ColBERT framework (Khattab & Zaharia, 2020) by treating image patches as visual tokens that interact with textual query tokens via MaxSim operations. Subsequent models built on this foundation – such as Llama-NemoRetriever-ColEmbed (Xu et al., 2025) and ColNomic-Embed-Multimodal (Team, 2025b) – have established state-of-the-art performance by capturing fine-grained interactions that are lost in single-vector representations. However, this performance comes at a significant cost: to achieve their results, these models rely on massively scaled representations, leading to substantial latency and storage overheads that can hinder practical deployment. This pressing trade-off between performance and efficiency motivates our work. Instead of pursuing ever-larger monolithic models, we investigate an alternative direction: enhancing these powerful vision-centric retrievers by fusing their signal with a complementary lightweight text-based encoder at test time.
@@
Numerous works apply hybrid retrieval in the context of combining dense semantic text retrieval with sparse lexical representations (Karpukhin et al., 2020; Kuzi et al., 2020; Luan et al., 2021; Chen et al., 2022). Bruch et al. (2023) conduct a theoretical and empirical analysis of the different ways to perform such dense-sparse fusions. Specifically, they compare RRF (Cormack et al., 2009) to score-based fusion, and analyze the sensitivity to the choice of tuned weights and normalizations. Hsu & Tzeng (2025) propose to set the score-fusion weights dynamically for each query at test time, based on (costly) feedback from an LLM judge. In contrast, GQR departs from these approaches by operating at test time on the representation level of the primary retriever.
@@
Prior work optimizes query representations during inference in text-only setups using pseudo-relevance feedback, often distilling from a cross-encoder re-ranker to a single-vector dense retriever (Yu et al., 2021; Sung et al., 2023; Gangi Reddy et al., 2025). Cross-encoders provide rich interactions, but incur substantial test-time cost. Here we replace the cross-encoder with a complementary bi-encoder (possibly of a different modality), which preserves low latency while still providing a strong guidance signal.
@@ 6 Conclusion
Through extensive experiments on the ViDoRe2 benchmarks, we demonstrated that GQR consistently improves retrieval performance across diverse model pairs, pushing ColPali-based retrievers to the latency–memory Pareto frontier. Our findings highlight that even weaker retrievers can provide valuable complementary guidance, underscoring the potential for resource-efficient retrieval systems in multimodal large-scale settings.
Limitations
GQR introduces a latency overhead at inference time, due to the additional retriever and iterative optimization process. We show that it is substantially more efficient than other base models with larger representations or cross-encoder rerankers, yet for applications with stringent real-time latency constraints, this additional cost should be taken into account. Secondly, GQR assumes access to a small, in-domain development set of queries for tuning its hyperparameters, and thus is not designed for a zero-shot setting. Finally, our focus in this work is empirical, testing the value of GQR in practical settings. Elucidating the theoretical foundations of GQR is an avenue for future work.
@@ Appendix A Background
Neural Information Retrieval represents a fundamental shift from traditional lexical matching methods like BM25 (Robertson et al., 1995), TF-IDF (Salton & Buckley, 1988), and other term-based approaches (Zhai & Lafferty, 2017). Unlike these sparse retrieval methods that rely on exact term matches and statistical properties, neural approaches learn dense semantic representations that capture conceptual similarity, enabling retrieval based on meaning rather than just shared vocabulary.
@@
have transformed information retrieval by learning semantic representations of queries and passages in a shared embedding space. These models typically produce a single dense vector representation for each passage and query, enabling efficient similarity computation through operations like cosine similarity or inner product. Early work by Karpukhin et al. (2020) introduced Dense Passage Retrieval (DPR), demonstrating that dense representations could outperform traditional sparse methods like BM25 for open-domain question answering. For query, $q$ and passage $p$, The similarity score $S$ is defined as:
@@
The BERT-based bi-encoder architecture, where queries and passages are independently encoded using separate or shared BERT models, became the foundational paradigm for neural retrieval. This framework has shaped the field, with numerous dense retrieval models building upon it (Reimers & Gurevych, 2019; Xiong et al., 2020; Qu et al., 2020; Izacard et al., 2021). The bi-encoder design enables pre-computation of passage embeddings, making real-time retrieval feasible at scale. Recently, advanced models such as Linq-Embed-Mistral (Choi et al., 2024) and Qwen3-Embedding (Zhang et al., 2025) extend this paradigm by leveraging larger language models as the underlying encoder to create more powerful dense representations.
@@
compute query-document similarity through more fine-grained interaction mechanisms. Rather than compressing all information into a single vector, these models preserve individual token embeddings for both queries and passages. ColBERT (Khattab & Zaharia, 2020) pioneered this approach with its MaxSim operation, which computes the maximum similarity between each query token and all passage tokens, then aggregates these scores. The MaxSim equation, as defined in ColBERT (see eq. 4)), finds for query token embedding $q_{i}$ the maximum similarity (dot product) with any passage token embedding $p_{j}$. These maximum scores are then summed up to get the final relevance score.
@@
This approach balances effectiveness and efficiency, as passage representations can still be pre-computed and indexed while the token-level matching captures finer details than single-vector approaches. Subsequent work improved efficiency and retrieval quality through various optimizations (Santhanam et al., 2021; 2022).
@@
forms the backbone of modern retrieval model optimization. These methods learn representations by pulling positive query-passage pairs closer while pushing negative pairs apart in the embedding space. The InfoNCE loss (Oord et al., 2018), which maximizes the similarity of positive pairs relative to negative ones through a softmax-like normalization, has become the dominant training objective across retrieval architectures.
@@
represent a different paradigm where query and passage are jointly encoded, enabling deep self-attention between their representations. Unlike bi-encoders that independently encode queries and passages, cross-encoders process query-passage pairs through a single model, allowing full attention across all tokens. This joint encoding is computationally expensive, making it impractical for first-stage retrieval over large corpora. However, cross-encoders excel as re-rankers, refining the top-K results from efficient first-stage retrievers. Examples include MonoBERT (Nogueira et al., 2019) for text retrieval and Chaffin & Lac (2024); Wasserman et al. (2025) for visual document retrieval.
@@ A.1 Visual Document Retrieval
were the dominant approach before the advent of end-to-end vision models. These pipelines convert visual documents into text through various techniques: traditional Optical Character Recognition (OCR) tools like docling (Auer et al., 2024) extract printed text, while Vision-Language Models (VLMs) can generate textual descriptions of visual elements such as charts, diagrams, and infographics. After verbalization, these methods apply standard text retrieval techniques to the extracted content. While verbalization-based approaches can leverage powerful text-only retrieval models, they inherently lose spatial relationships and visual context during the text extraction process.
@@
introduced a new approach to visual document retrieval by directly encoding document images without intermediate text extraction. These VLM-based embedding models transform pre-trained generative Vision-Language Models (such as PaliGemma, Beyer et al., 2024 and Qwen-VL, Wang et al., 2024) into multi-vector embedding models optimized for retrieval.
ColPali (Faysse et al., 2025) was the first model to introduce this approach, building upon PaliGemma (Beyer et al., 2024) and adapting the late-interaction framework to vision-language models by treating image patches as visual tokens that interact with textual query tokens through MaxSim operations. ColPali provides native text-query support due to its VLM-based design. Queries remain text, are encoded by the model’s language tower, and are matched directly against visual page/passage tokens, eliminating any OCR at query time. Training is contrastive, typically InfoNCE-style with in-batch/hard negatives to align query tokens with relevant visual tokens. This architecture preserves spatial layout information and visual features that verbalization-based methods discard. Following ColPali’s success, subsequent models have adopted and extended this ColPali paradigm by leveraging different base VLMs: Llama-NemoRetriever-ColEmbed (Xu et al., 2025), Colnomic-Embed-Multimodal (Team, 2025b), Granite-Vision-Embedding (Team, 2025a) and Jina-Embeddings-v4 (Günther et al., 2025).

### v2

@@ 3.3 Comparison to hybrid retrieval pipelines
Table 3: Percentage gain in $\mathrm{NDCG}@5$ of hybrid retrieval over the primary retriever on ViDoRe 2. Each cell depicts the average gain over 99 retriever pairs (33 multimodal base retrievers and $3$ text retrievers).
Method Avg Std Per-Subset Gain
Biomed Lectures Economics ESG Human ESG Full
@@
Table 3 depicts the aggregated performance of the different hybrid retrieval methods on ViDoRe 2, presented as the average percentage gain relative to the base retriever over all pairs (Table 6 in the appendix lists the average absolute values). The Std column captures the standard deviation of performance across the pairs for each method. It shows that the ranking aggregation methods (RRF, Average Ranking) generally lead to a deterioration in performance. All aggregation methods benefit from parameter tuning, with tuned score aggregation methods achieving consistent gains over the base retriever. Table 5 in Appendix F shows the results in zero-shot settings on ViDoRe 3 (see Figure 7 for how the baseline zero-shot configuration is chosen).
Across the benchmarks, GQR outperforms all other hybrid retrieval variants, with the largest average gains over the base multimodal retriever, while also emerging as the most stable positive method with a low standard deviation across pairs.
3.4 Results - Efficiency
@@
To characterize the test-time costs of GQR and ColPali-based retrievers, we ran latency measurements on a single NVIDIA A100 GPU, measuring document retrieval for randomly sampled queries from each ViDoRe 2 subset (averaging across $100$ runs). Figure 3 depicts the quality-latency trade-offs of the different retrievers, where each point on the graph represents either a base retriever or a retriever with GQR (Appendix F shows full values). The plot illustrates the substantial costs of the strongest base model – Llama-Nemo (orange square, right) – which attains $\mathrm{NDCG}@5=62.9$ at a cost of $2{,}591\,\text{ms}$ per query. Notably, our Colnomic GQR hybrid, with Linq as the refinement model (green diamond, left), reaches $\mathrm{NDCG}@5=62.7$ at $181\,\text{ms}$ ($\approx\!14\times$ faster), and the Colnomic hybrid with Jina (green circle) attains $\mathrm{NDCG}@5=63.0$ at $350\,\text{ms}$ ($\approx\!7\times$ faster), surpassing Nemo. Across base models, applying GQR increases latency by small relative measures with large gains in performance, shifting the Pareto frontier left and upwards. Appendix F shows that GQR is also on the index-storage Pareto frontier.
Figure 3: Latency–quality tradeoff in online querying. The $x$ axis is runtime in milliseconds for a single query, on a log scale, and the $y$ axis is the average evaluation score (NDCG@5). Empty squares indicating the primary retriever alone (without applying GQR).
@@ 4 Analysis
4.1 Hyperparameter investigation
GQR relies on two hyperparameters - the learning rate $\alpha$ and the number of optimization steps $T$. We evaluate performance over the Cartesian product $\{1\times 10^{-5},\,5\times 10^{-5},\,10^{-4},\,5\times 10^{-4},\,10^{-3},\,5\times 10^{-3}\}\times\{5,10,15,\ldots,90,95,100\}$ of $\alpha$ and $T$. We run this analysis over 6 model pairs, excluding Llama-NemoRetriever-Colembed-3B due to its computational overhead.
Figure 4: Hyperparameter sweep over GQR’s learning rate $\alpha$ and optimization steps $T$, averaged over six model pairs on ViDoRe 2.
Figure 5: Online latency breakdown of GQR for $T=25$ and $T=50$.
Figure 5 depicts GQR performance on ViDoRe 2 for each combination of $\alpha$ and $T$, averaged over the $6$ model pairs. It shows that high learning rates ($10^{-3},\,5\times 10^{-4}$) are compatible with small $T$ values (1010-$25$), where beyond that performance starts decreasing. Medium learning rates ($10^{-4},\,5\times 10^{-5}$), show consistent improvement and stable optimization, yet require more steps (7575-$100$) to reach higher performance. The smallest ($10^{-5}$) and largest ($5\times 10^{-3}$) learning rates are suboptimal, where the latter even results in performance degradation relative to the primary retriever.
The results capture a tradeoff between latency and stability. Higher learning rates can provide a performance boost faster, but might deteriorate quickly past a certain $T$ value, while medium learning rates are more stable yet require more online optimization steps. Overall, three configurations emerge as the most promising for GQR: (1) $\alpha=10^{-3}$ with $5<T<15$, (2) $\alpha=5\times 10^{-4}$ with $15<T<30$, and (3) $\alpha=10^{-4}$ with $50<T<75$. We leave it to future work to explore methods for optimizing this tradeoff, potentially by incorporating learning-rate scheduling procedures.
4.2 Online latency breakdown
Figure 5 describes the online latency breakdown of GQR, for $T=25$ and $T=50$. The primary model used is Colnomic-Embed-Multimodal-7B and the complementary is Linq-Embed-Mistral. Latencies are collected with an A100 gpu over the ViDoRe 2 benchmark.
The test-time optimization accounts for 50 ms of a 192 ms total latency for $T=25$ and 99 ms of a 241 ms total for $T=50$, averaging roughly 2 ms per optimization step. The procedure is shown here in a sequential form, although with sufficient GPU memory the query encoding and index search for the two retrievers can be parallelized, reducing overall latency.
@@ 4.3 Reranker Comparison
Cross-encoder rerankers are widely used in retrieval pipelines, applying full query to document interaction via self-attention to the top-K documents. Similarly to GQR, rerankers operate at test time on a candidate pool returned by a bi-encoder, and thus provide a natural point of comparison. We evaluate GQR against lightonai/MonoQwen2-VL-v0.1, an open-weights multimodal reranker (MonoQwen). Figure 6 illustrates the latency-performance characteristics of GQR against reranking the top-$5$ (on the left) and the top-$10$ (on the right) candidates returned by each retriever. GQR is run with a default $K=10$ configuration and with Linq-Embed as the refinement model. The left plot illustrates that GQR outperforms a top-$5$ reranking pipeline on both the latency and performance axes. Against the top-$10$ reranking pipelines (right), GQR achieves close performance while being $21\times$ faster for Colnomic, $16\times$ faster for Jina (vision), and $2\times$ faster for Nemo. Across both comparisons, GQR remains on the Pareto front and indicates attractive latency performance trade-offs. We report the full results in Table 11.
@@ 4.4 GQR Design Choices
Prior works on test-time query optimization often run a second index search with the optimized query $z^{(T)}$, aiming to retrieve documents beyond the initial top-$K$ and increase recall (sung-etal-2023-optimizing; 10.1145/3726302.3730160). This extra pass adds latency as it repeats index traversal and candidate generation. We test this modification to GQR where we perform an additional search with the optimized query over the full index, and observe no improvement in performance (Table 16). This suggests that the effects of GQR in our setting are largely confined to the original pool $\mathcal{C}(q)$ of candidate documents.
@@
We conduct additional experiments on GQR design choices in Appendix E.
@@ 5 Related Works
Recent advances in visual document retrieval have been dominated by late-interaction, multi-vector architectures. This paradigm, first introduced to the multimodal domain by ColPali (faysse2025colpali), adapts the ColBERT framework (khattab2020colbert) by treating image patches as visual tokens that interact with textual query tokens via MaxSim operations. Subsequent models built on this foundation – such as Llama-NemoRetriever-ColEmbed (xu2025llama_nemo) and ColNomic-Embed-Multimodal (nomicembedmultimodal2025) – have established state-of-the-art performance by capturing fine-grained interactions that are lost in single-vector representations. However, this performance comes at a significant cost: to achieve their results, these models rely on massively scaled representations, leading to substantial latency and storage overheads that can hinder practical deployment. This pressing trade-off between performance and efficiency motivates our work. Instead of pursuing ever-larger monolithic models, we investigate an alternative direction: enhancing these powerful vision-centric retrievers by fusing their signal with a complementary lightweight text-based encoder at test time.
@@
Numerous works apply hybrid retrieval in the context of combining dense semantic text retrieval with sparse lexical representations (karpukhin-etal-2020-dense; kuzi2020leveraging; luan-etal-2021-sparse; chen2022out). bruch2023analysis conduct a theoretical and empirical analysis of the different ways to perform such dense-sparse fusions. Specifically, they compare RRF (36196) to score-based fusion, and analyze the sensitivity to the choice of tuned weights and normalizations. hsu2025dat propose to set the score-fusion weights dynamically for each query at test time, based on (costly) feedback from an LLM judge. In contrast, GQR departs from these approaches by operating at test time on the representation level of the primary retriever.
@@
Prior work optimizes query representations during inference in text-only setups using pseudo-relevance feedback, often distilling from a cross-encoder re-ranker to a single-vector dense retriever (yu2021improvingqueryrepresentationsdense; sung-etal-2023-optimizing; 10.1145/3726302.3730160). Cross-encoders provide rich interactions, but incur substantial test-time cost. Here we replace the cross-encoder with a complementary bi-encoder (possibly of a different modality), which preserves low latency while still providing a strong guidance signal.
@@ 6 Conclusion
Through extensive experiments on the ViDoRe benchmarks, we demonstrated that GQR consistently improves retrieval performance across diverse model pairs, pushing ColPali-based retrievers to the latency–memory Pareto frontier. Our findings highlight that even weaker retrievers can provide valuable complementary guidance, underscoring the potential for resource-efficient retrieval systems in multimodal large-scale settings.
@@ Appendix A Background
Neural Information Retrieval represents a fundamental shift from traditional lexical matching methods like BM25 (robertson1995okapi_bm25), TF-IDF (salton1988term_tf_idf), and other term-based approaches (zhai2017study_term_based). Unlike these sparse retrieval methods that rely on exact term matches and statistical properties, neural approaches learn dense semantic representations that capture conceptual similarity, enabling retrieval based on meaning rather than just shared vocabulary.
@@
have transformed information retrieval by learning semantic representations of queries and passages in a shared embedding space. These models typically produce a single dense vector representation for each passage and query, enabling efficient similarity computation through operations like cosine similarity or inner product. Early work by karpukhin-etal-2020-dense introduced Dense Passage Retrieval (DPR), demonstrating that dense representations could outperform traditional sparse methods like BM25 for open-domain question answering. For query, $q$ and passage $p$, The similarity score $S$ is defined as:
@@
The BERT-based bi-encoder architecture, where queries and passages are independently encoded using separate or shared BERT models, became the foundational paradigm for neural retrieval. This framework has shaped the field, with numerous dense retrieval models building upon it (reimers2019sentence_bert; xiong2020approximate; qu2020rocketqa; izacard2021unsupervised). The bi-encoder design enables pre-computation of passage embeddings, making real-time retrieval feasible at scale. Recently, advanced models such as Linq-Embed-Mistral (choi2024linq) and Qwen3-Embedding (qwen3embedding) extend this paradigm by leveraging larger language models as the underlying encoder to create more powerful dense representations.
@@
compute query-document similarity through more fine-grained interaction mechanisms. Rather than compressing all information into a single vector, these models preserve individual token embeddings for both queries and passages. ColBERT (khattab2020colbert) pioneered this approach with its MaxSim operation, which computes the maximum similarity between each query token and all passage tokens, then aggregates these scores. The MaxSim equation, as defined in ColBERT (see eq. 4)), finds for query token embedding $q_{i}$ the maximum similarity (dot product) with any passage token embedding $p_{j}$. These maximum scores are then summed up to get the final relevance score.
@@
This approach balances effectiveness and efficiency, as passage representations can still be pre-computed and indexed while the token-level matching captures finer details than single-vector approaches. Subsequent work improved efficiency and retrieval quality through various optimizations (santhanam2021colbertv2; santhanam2022plaid).
@@
forms the backbone of modern retrieval model optimization. These methods learn representations by pulling positive query-passage pairs closer while pushing negative pairs apart in the embedding space. The InfoNCE loss (oord2018representation_infonce), which maximizes the similarity of positive pairs relative to negative ones through a softmax-like normalization, has become the dominant training objective across retrieval architectures.
@@
represent a different paradigm where query and passage are jointly encoded, enabling deep self-attention between their representations. Unlike bi-encoders that independently encode queries and passages, cross-encoders process query-passage pairs through a single model, allowing full attention across all tokens. This joint encoding is computationally expensive, making it impractical for first-stage retrieval over large corpora. However, cross-encoders excel as re-rankers, refining the top-K results from efficient first-stage retrievers. Examples include MonoBERT (nogueira2019multi_mono_bert) for text retrieval and MonoQwen; wasserman2025docrerank for visual document retrieval.
@@ A.1 Visual Document Retrieval
were the dominant approach before the advent of end-to-end vision models. These pipelines convert visual documents into text through various techniques: traditional Optical Character Recognition (OCR) tools like docling (auer2024docling) extract printed text, while Vision-Language Models (VLMs) can generate textual descriptions of visual elements such as charts, diagrams, and infographics. After verbalization, these methods apply standard text retrieval techniques to the extracted content. While verbalization-based approaches can leverage powerful text-only retrieval models, they inherently lose spatial relationships and visual context during the text extraction process.
@@
introduced a new approach to visual document retrieval by directly encoding document images without intermediate text extraction. These VLM-based embedding models transform pre-trained generative Vision-Language Models (such as PaliGemma, beyer2024paligemma and Qwen-VL, wang2024qwen2) into multi-vector embedding models optimized for retrieval.
ColPali (faysse2025colpali) was the first model to introduce this approach, building upon PaliGemma (beyer2024paligemma) and adapting the late-interaction framework to vision-language models by treating image patches as visual tokens that interact with textual query tokens through MaxSim operations. ColPali provides native text-query support due to its VLM-based design. Queries remain text, are encoded by the model’s language tower, and are matched directly against visual page/passage tokens, eliminating any OCR at query time. Training is contrastive, typically InfoNCE-style with in-batch/hard negatives to align query tokens with relevant visual tokens. This architecture preserves spatial layout information and visual features that verbalization-based methods discard. Following ColPali’s success, subsequent models have adopted and extended this ColPali paradigm by leveraging different base VLMs: Llama-NemoRetriever-ColEmbed (xu2025llama_nemo), Colnomic-Embed-Multimodal (nomicembedmultimodal2025), Granite-Vision-Embedding (granite_vision_embedding) and Jina-Embeddings-v4 (gunther2025jina_v4).